OpenStudy (anonymous):

Please Help! Write the equation of an ellipse with a major axis of length 8 and co-vertices (0, 3) and (0,-3).

OpenStudy (amistre64):

how far did you get?

OpenStudy (amistre64):



OpenStudy (amistre64):

\[(\frac x4)^2+(\frac y3)^2=1\]

OpenStudy (anonymous):

I ended up getting x^2/8+y^2/sqrt73=1 but I don't think it is right

OpenStudy (amistre64):

yeah, yours is not right

OpenStudy (anonymous):

How did you get your answer then?

OpenStudy (amistre64):

take the standard equation of an ellipse centered at the point (h,k) \[(\frac{x-h}{a})^2+(\frac{y-k}{b})^2=1\] a and b are measures for the x and y from the CENTER we are given that the total distance for one of them is 8, half of 8 is: 4 we are given that the other measure spans from (0,3) to (0,-3); this tells us 2 things. The center is at (0,0) and the half distance is 3 on the y parts; which leaves the 4 to be on the x part

OpenStudy (amistre64):

\[(\frac{x-h}{a})^2+(\frac{y-k}{b})^2=1\] \[(\frac{x-0}{a})^2+(\frac{y-0}{b})^2=1\] \[(\frac{x}{a})^2+(\frac{y}{3})^2=1\] \[(\frac{x}{4})^2+(\frac{y}{3})^2=1\]

OpenStudy (amistre64):